# Physical modelling of global macrosystems evolution

### v1

@@ III.2 World Population and GDP
The growth of world population size over the years surpasses exponential rates, as widely recognized in prior literature Kremer (1993); Kapitza (1996). The seminal paper from 1960, ”Doomsday: Friday, 13 November, A.D. 2026” Von Foerster et al. (1960), was among the first to identify this phenomenon. The authors observed that demographic growth from the year 1 CE to 1958 CE followed a hyperbolic pattern:
@@
with approximate constants $C\approx 2.15\cdot 10^{5}$ and $t^{*}\approx 2026$. The value $t^{*}$ represents the mathematical singularity of the hyperbolic curve, which was used to speculate about future population growth. This hyperbolic pattern has been verified from earlier periods Kremer (1993); Kapitza (1996) up to a million years ago and continues to be relevant for some years beyond 1958 Korotayev (2020).
@@
Typically, these datasets are interpreted using dynamic models that consider the tradeoff between mortality and fecundity Lee (1980) without explicit accounting for individual interaction. In contrast, the combinatorial approach based on TAP has been employed to explain world population growth Koppl et al. (2018), although it does not focus on the specific types of interactions involved. Here, we show that by incorporating particular interaction terms into the combinatorial model, we can reproduce and predict the observed hyperbolic growth.
@@
Starting with the general TAP equation 1, we select appropriate values for the parameters $\alpha_{i}$. Demographic growth is driven by reproductive processes, which can be modeled as pair interactions. We then hypothesize that the driving term is for $i=2$, corresponding to the number of pairs in the system:
$\dfrac{dN_{t}}{dt}=\alpha^{2}\binom{N_{t}}{2}=\alpha^{2}\dfrac{N_{t}(N_{t}-1)}{2}\approx\gamma N_{t}^{2}$ (7)
@@
Given $N_{t}\gg 1$, we can neglect the linear term, focusing on the quadratic term and redefining $\gamma=\alpha^{2}/2$ as a birth rate. Notice that we do not consider direct mortality effects, so $\mu=0$ in equation 1.
@@
The solution to equation 7 produces the hyperbolic curve (details are reported in Appendix B):
$N_{t}\approx\dfrac{1}{N_{0}^{-1}-\gamma(t-t_{0})}$ (8)
@@
This result aligns with previous analyses Korotayev (2020). It is important to note that this solution is obtained using the continuous approximation of the model, while a discrete model would not exhibit any singularity. Results of the fit of the macrodevelopment parameter $\gamma$ are reported in Figure 2 and in Appendix C.
Figure 3: Evolution of the number of patents and technological codes over time. Left. Number of patents over time, shown on a log-linear scale. There is a rapid growth phase until approximately 1860, after which the trend becomes exponential. The solid line represents an exponential fit of the data starting from 1900. The inset shows the growth of the number of distinct technological codes in time on a linear scale. The solid line, following the empirical trend, highlights the linear growth of technological codes. Right. Number of patents as a function of technological codes, presented on a log-linear scale, which highlights the exponential growth of patents. As discussed in Appendix C, the solid line represents the fit. The inset shows the number of technological codes versus the number of patents. In this case, the trend is reversed. The linear-log scale highlights the logarithmic growth of technological codes. The solid line represents the model fit.
@@
Furthermore, some works Korotayev and LePoire (2020); Korotayev (2020) mistakenly equate the hyperbolic patterns of demographic growth with those of Canonical Milestones, suggesting they are generated by the same process. However, these situations are distinct, as demonstrated by the different models used to fit the data. For Canonical Milestones ($M_{t}$), the hyperbolic pattern pertains to the rate, described by the equation 5, which involves all combinatorial terms. In contrast, for the world population, the hyperbolic pattern emerges directly in $N_{t}$ (Eq. 8), where only pair interactions are considered.
@@
Finally, we extend our analysis to GDP per capita, reaffirming previous findings within the new combinatorial framework. Many researchers have postulated a relationship between GDP growth and world population Malthus (2023); Acemoglu (2008). According to Malthusian macroeconomic principles, the GDP $K_{t}$ grows at a rate proportional to the size of the population $N_{t}$. We can represent this growth rate with a combinatorial term proportional to $K_{t}N_{t}$ Korotayev (2009). Additionally, empirical data show that GDP’s behavior over the years closely mirrors the square of the population $N_{t}\approx\sqrt{K_{t}}$ Korotayev (2005). Using these observations, the GDP growth rate can be expressed as:
$\dfrac{dK_{t}}{dt}=\delta K_{t}N_{t}\approx\delta K_{t}^{3/2}$ (9)
@@
where $\delta$ is the macrodevelopment parameter. The solution is a hyperbolic curve with an exponent of 2 (analytical details are given in Appendix B):
$K_{t}\approx\dfrac{1}{\left(K_{0}^{-1/2}-\delta t\right)^{2}}$ (10)
@@
This result matches the expression fitted to the data in previous works Korotayev (2009). Once again, this hyperbolic behavior is explained within the combinatorial growth framework. Results of the fit of the parameter $\delta$ are reported in Figure 2 and Appendix C.
@@ III.3 Technological Innovations: the Case of US Patents
Technological innovations, as demonstrated through patenting activities, provide an effective example of combinatorial growth. Patents are characterized by a set of codes, which reveal the underlying technological capabilities used in the invention. As a result, patents naturally emerge as combinations of technological codes. However, this system differs significantly from those previously explored, as the number of patents displays an exponential, rather than hyperbolic, increase over time Youn et al. (2015).
Due to this exponential behavior, previous frameworks have either omitted the example of US patents or approached it differently. In some instances, the TAP equation was employed to explain the distribution of descendants of any given patent, with minimal consideration for their temporal evolution Koppl et al. (2023). In another study Youn et al. (2015), the authors conducted a detailed statistical analysis of the US patents dataset, focusing on the combinatorial processes underlying the creation of new technologies. However, no model was proposed to reproduce the observed dynamics.
It is easy to show that our framework can also describe the dynamics of patents and reproduce the essential features of the data. The left panel of Figure 3 illustrates the exponential growth of the number of patents over time when plotted on a log-linear scale. Conversely, the behavior of technological codes is linear rather than exponential, as shown in the inset. As shown in the right panel of Figure 3, plotting the number of patents relative to the number of distinct technological codes reveals an exponential relationship.
To build the model, we exploit the idea that patents are created through the combination of technological codes. Patents can theoretically involve any number of codes, with larger combinations presenting more challenges to be realized. Therefore, we consider all terms within equation 1 to establish an exponential model, similar to the approach outlined in Section III.1:
@@
with $\beta=\ln(1+\alpha)$, where $\alpha$ is the usual macrodevelopment rate of the system. Here, $C$ represents the number of technological codes, while $P_{C}$ denotes the number of patents associated with a given code $C$. In this model, we treat the number of technological codes as the timescale, $C$. Consequently, the solution of equation 11 is:
$P_{C}\propto\dfrac{1}{\beta}e^{\beta C},\quad C_{P}\propto\dfrac{1}{\beta}\log\beta P$ (12)
This equation replicates the observed trends of patents and technological codes, as shown in the right panel of Figure 3. We also present the specific curve obtained by fitting the data. Further details regarding the coefficients used in this model are provided in Appendix C.
Moreover, the relationships outlined in equation 12 enable us to recover the temporal growth of these two quantities, as shown in the left panel of Figure 3. In regions where $C(t)$ approximates linearity, an exponential growth pattern emerges for patents, such that $P(t)\propto e^{t}$. It is important to stress that, in this case, the model does not produce any singularity, even if the underlying mechanism is analogous to that applied to other systems. This showcases the versatility of the TAP equation and the theory of combinatorial innovation to describe a wide variety of different systems with similar but yet different underlying dynamics.
@@ IV Discussion
The mathematical singularity arising from the hyperbolic behaviors described in Section III has been subject to various interpretations. The concept of “Singularity” was already mentioned in the work of Von Foerster et al. Von Foerster et al. (1960), who provocatively identified a specific date for the so-called ”Doomsday,” projected to occur on 13 November 2026. More recently, scholars offered two main interpretations of this concept. The first, as suggested by Von Foerster et al. (1960), argues that such a singularity indicates that rapid growth is unsustainable in real finite systems. This view implies that a slowdown in growth must occur before any potential explosive scenario Korotayev (2020); Korotayev and LePoire (2020); Modis (2013). In this interpretation, the mathematical divergence can be seen as a potential indicator of this deceleration getting closer. Indeed, in many systems, it is reasonable to expect a deceleration due to the presence of limiting factors such as the finite availability of resources. Nonetheless, our mathematical analysis reveals that inferring this slowdown from the data may be very tricky. First, the discrete versions of the models, which represent the actual dynamics of real-world systems, do not exhibit any singularity. The discrete models can be indeed used to forecast growth beyond this theoretical limit, as detailed in Appendix C. Secondly, our models do not explicitly account for external limiting factors, which are the true drivers of any changes in the system. As a result, these simple models are not able to predict any shift in the system’s evolution. This makes the singularity value practically unrelated to any constraints or deceleration.
@@
Conversely, some authors, particularly in the Big History field, view the singularity as an indicator of upcoming transitions or technological shifts Kurzweil (2005); Panov (2020); Nazaretyan (2015). For instance, the concept of Threshold 9 in the field of Big History Nazaretyan (2016) suggests that humanity is approaching a profound global paradigm shift (the threshold) that will significantly impact both technology and evolution Kurzweil (2005). Some authors even speculate about post-singularity civilizations Panov (2011). This second interpretation often relies on extrapolations from data without direct evidence that current growth rates will continue indefinitely, positioning such analyses more within the realm of philosophical speculation, though they inspire fascinating and stimulating discussions. From a technical perspective, attempts to extrapolate precise years from the data can lead to peculiar results. For example, in Von Foerster et al. (1960), the authors calculated the value $N_{t}=1$ from the world population curve and interpreted it as ”Adam’s birth,” resulting in a date nearly 200 billion years in the past. Similarly, some predictions appear strikingly accurate but are essentially coincidental, such as the singularity in GDP per capita growth estimated for 2005 Korotayev (2009), which is closely aligned with the 2008 financial crisis.
@@
Finally, it is important to stress that estimating a reliable singularity value is almost impossible due to the high value of statistical error. As noted in Von Foerster et al. (1960), the year obtained from fitting data can vary substantially depending on the choice of the last observation. For instance, if Charlemagne, with data until 800 CE, had predicted a singularity, he would have projected it 300 years into the future, while Napoleon, using more recent data, would have predicted it just 30 years ahead. In the very same way, our estimates of singularities, presented in Appendix C, differ from previous analyses in the literature due to the updated data. In certain cases, large time intervals considered in the analysis result in wide confidence bounds for the numerical value of the singularity, making it even more difficult to obtain a reliable estimate.
@@ V Conclusions
This study investigates the dynamics of growth and innovation, offering a unified approach for modeling and understanding it. Innovation dynamics often exhibit hyperbolic trends, which have been studied and interpreted in various ways in the literature. However, a coherent framework that models these processes from a physical and mathematical perspective has been lacking. By leveraging a combinatorial approach based on the Theory of Adjacent Possible (TAP), we provide a comprehensive theoretical foundation to interpret these phenomena. We investigated several distinct scenarios, demonstrating how our framework can successfully replicate the main characteristics of growth trends. We analyzed four key cases: biological milestones (Canonical Milestones and Biological Phase Transitions), world population, GDP growth, and technological innovations as seen through U.S. patent data. In each case, the models allow us to estimate the macrodevelopment parameter from the data, enabling us to quantify the system’s growth rate.
This paper demonstrates how a combinatorial framework can offer a coherent interpretation of growth and innovation processes across diverse domains. Future research should explore applying these models to other complex systems, such as biological or evolutionary processes, where similar growth dynamics are observed. The main limitation of our approach is that it does not consider any external factor acting on the system, which is relevant in most cases. Given the simplicity and interpretability of our approach, future work could focus on incorporating external constraints, such as resource limitations, to produce more accurate descriptions and forecasts of system behavior. Another promising direction is investigating the connection between this framework and the triggering mechanisms described in other discovery processes Tria et al. (2014); Di Bona et al. (2023); Bellina et al. (2024). In this view, the emergence of innovations is explained through exploring a network, where each novelty opens up the possibility for further developments. Establishing a connection between these frameworks could further advance our understanding of innovation processes and achieve a comprehensive theory of innovation.

### v2

@@ III.2 World Population and GDP
Figure 3: Evolution of the number of patents and technological codes over time. Left. Number of patents over time, shown on a log-linear scale. There is a rapid growth phase until approximately 1860, after which the trend becomes exponential. The solid line represents an exponential fit of the data starting from 1900. Time represents the number of years in the Common Era (CE). Center. Number of distinct technological codes as function of time on a linear scale. The solid line, following the empirical trend, highlights the linear growth of technological codes observed up to $1980$. Right. Number of patents as a function of technological codes, presented on a log-linear scale, which highlights the exponential growth of patents. As discussed in Appendix C, the solid line represents the fit.
The growth of world population size over the years surpasses exponential rates, as widely recognised in prior literature kremer1993population ; kapitza1996phenomenological . The seminal paper from 1960, ”Doomsday: Friday, 13 November, A.D. 2026” von1960doomsday , was among the first to identify this phenomenon. The authors observed that demographic growth from the year 1 CE to 1958 CE followed a hyperbolic pattern:
@@
with approximate constants $C\approx 2.15\cdot 10^{5}\ [y]$ and $t^{*}\approx 2026\ [y]$. The value $t^{*}$ represents the mathematical singularity of the hyperbolic curve, which was used to speculate about future population growth. This hyperbolic pattern has been verified from earlier periods kremer1993population ; kapitza1996phenomenological up to a million years ago and continues to be relevant for some years beyond 1958 korotayev2020twenty .
@@
Typically, these datasets are interpreted using dynamic models that consider the tradeoff between mortality and fecundity lee1980aiming without explicit accounting for individual interaction. In contrast, the combinatorial approach based on TAP has been employed to explain world population growth koppl2018simple , although it does not focus on the specific types of interactions involved. Here, we show that by incorporating particular interaction terms into the combinatorial model, we can reproduce and predict the observed hyperbolic growth.
@@
Starting with the general TAP equation (1), we select appropriate values for the parameters $\alpha_{i}$. Demographic growth is driven by reproductive processes, which can be modelled as pair interactions. We then hypothesize that the driving term is for $i=2$, corresponding to the number of pairs in the system:
$\dfrac{dN_{t}}{dt}=\overline{\nu}\alpha^{2}\binom{N_{t}}{2}=\overline{\nu}\alpha^{2}\dfrac{N_{t}(N_{t}-1)}{2}\approx\overline{\nu}\gamma N_{t}^{2}=\nu_{\gamma}N_{t}^{2}$ (11)
@@
Given $N_{t}\gg 1$, we can neglect the linear term, focusing on the quadratic term and redefining $\gamma=\alpha^{2}/2$ as a birth rate. Notice that we do not consider direct mortality effects, so $\mu=0$ in Eq. (1). Additionally, the usual dimensional factor $\overline{\nu}$ is reabsorbed into the parameter $\nu_{\gamma}=\overline{\nu}\gamma$, which now carries the dimension of a rate.
@@
The solution to equation (11) produces the hyperbolic curve (details are reported in Appendix B):
$N_{t}\approx\dfrac{1}{N_{0}^{-1}-\nu_{\gamma}(t-t_{0})}$ (12)
@@
This result aligns with previous analyses korotayev2020twenty . It is important to note that this solution is obtained using the continuous approximation of the model, while a discrete model would not exhibit any singularity. Results of the fit of the macrodevelopment parameter $\gamma$ are reported in Figure 2 and in Appendix C. As in the previous case, the time scale is shifted by convention to represent the number of years before $2020$ CE.
@@
Furthermore, some works korotayev202021st ; korotayev2020twenty mistakenly equate the hyperbolic patterns of demographic growth with those of Canonical Milestones, suggesting they are generated by the same process. However, these situations are distinct, as demonstrated by the different models used to fit the data. For Canonical Milestones ($M_{t}$), the hyperbolic pattern pertains to the rate, described by the equation (8), which involves all combinatorial terms. In contrast, for the world population, the hyperbolic pattern emerges directly in $N_{t}$ (Eq. (12)), where only pair interactions are considered.
@@
Finally, we extend our analysis to GDP per capita, reaffirming previous findings within the new combinatorial framework. Many researchers have postulated a relationship between GDP growth and world population malthus2023essay ; acemoglu2008introduction . According to Malthusian macroeconomic principles, the GDP $K_{t}$ grows at a rate proportional to the size of the population $N_{t}$. We can represent this growth rate with a combinatorial term proportional to $K_{t}N_{t}$ korotayev2009compact . Additionally, empirical data show that GDP’s behaviour over the years closely mirrors the square of the population $N_{t}\approx\sqrt{K_{t}}$ korotayev2005compact . Using these observations, the GDP growth rate can be expressed as:
$\dfrac{dK_{t}}{dt}=\overline{\nu}\delta K_{t}N_{t}\approx\nu_{\delta}K_{t}^{3/2}$ (13)
@@
where $\nu_{\delta}$ is the macrodevelopment parameter, which has the dimension of a rate. The solution is a hyperbolic curve with an exponent of 2 (analytical details are given in Appendix B):
$K_{t}\approx\dfrac{1}{\left(K_{0}^{-1/2}-\nu_{\delta}(t-t_{0})\right)^{2}}$ (14)
@@
This result matches the expression fitted to the data in previous works korotayev2009compact . Once again, this hyperbolic behaviour is explained within the combinatorial growth framework. Results of the fit of the parameter $\delta$ are reported in Figure 2 and Appendix C.
@@ III.3 Technological Innovations: the Case of US Patents
Technological innovations, as demonstrated through patenting activities, provide an effective example of combinatorial growth. Patents are characterised by a set of codes, which reveal the underlying technological capabilities used in the invention. As a result, patents naturally emerge as combinations of technological codes. However, this system differs significantly from those previously explored, as the number of patents displays an exponential, rather than hyperbolic, increase over time youn2015invention .
Due to this exponential behaviour, previous frameworks have approached US patents differently youn2015invention . Specifically, the TAP equation has been used to model the temporal evolution of patents, as well as other aspects such as the distribution of descendants sole2016singularities ; koppl2023explaining . However, the absence of a singularity in the case of US patents requires introducing an additional aging process into the model. This process slows down the hyperbolic growth, leading to the emergence of an exponential behaviour. Here, we demonstrate that our framework captures the dynamics of patents, reproducing the essential features of the data directly from the TAP equation.
The left panel of Figure 3 illustrates the exponential growth of the number of patents over time when plotted on a log-linear scale. Conversely, the behaviour of technological codes is linear rather than exponential. As shown in the right panel of Figure 3, plotting the number of patents relative to the number of distinct technological codes reveals an exponential relationship.
Notice that, unlike previous cases, here patents and technological codes do not share the same nature. They represent, indeed, two different conceptual categories. For this reason, we exploit the idea that patents are created through the combination of technological codes, and we model the evolution of patents with respect to these codes. Patents can theoretically involve any number of codes, with larger combinations presenting more challenges to be realised. Therefore, we consider all terms within equation (1) to establish an exponential model, similar to the approach outlined in Section III.1:
@@
with $\beta=\ln(1+\alpha)$, where $\alpha$ is the usual macrodevelopment rate of the system. For simplicity, we treat all quantities (patents and technological codes) as dimensionless, as time is not explicitly present in this model. As a consequence, in this case, there is no need to explicitly consider any dimensional factor. Here, $C$ represents the number of technological codes, while $P_{C}$ denotes the number of patents when the number of technological codes is $C$. In this model, we treat the number of technological codes as a time scale $C$. Consequently, the solution of equation (15) is as follows:
$P_{C}\propto\dfrac{1}{\beta}e^{\beta C},\quad C_{P}\propto\dfrac{1}{\beta}\ln\beta P$ (16)
This equation replicates the trends observed for patents and technological codes, as shown in the right panel of Figure 3. We also present the specific curve obtained by fitting the data. More details on the coefficients used in this model are provided in Appendix C.
Moreover, the relationships described in Eq. (16) enable us to recover the temporal growth of the number of patents, as shown in the left panel of Figure 3. In regions where $C(t)$ approximates linearity, an exponential growth pattern emerges for patents such that $P(t)\propto e^{t}$. It should be noted, however, that the exponential relationship between the two quantities described by equation (16) remains valid regardless of the specific temporal behaviour of $C(t)$ or $P(t)$. It is important to stress that, in this case, the model does not produce any singularity, even if the underlying mechanism is analogous to that applied to other systems. This showcases the versatility of the TAP equation and the theory of combinatorial innovation to describe a wide variety of different systems with similar but yet different underlying dynamics.
@@ IV Discussion
The mathematical singularity arising from the hyperbolic behaviours described in Section III has been subject to various interpretations. The concept of “Singularity” was already mentioned in the work of Von Foerster et al. von1960doomsday , who provocatively identified a specific date for the so-called ”Doomsday,” projected to occur on 13 November 2026. More recently, scholars offered two main interpretations of this concept. The first, as suggested by von1960doomsday , argues that such a singularity indicates that rapid growth is unsustainable in real finite systems. As discussed in koppl2023explaining , a singularity ”would be a regime change. It would mark the point at which our model no longer applies”, highlighting its conceptual importance as a transition point, beyond which standard modelling may fail. This view implies that a slowdown in growth must occur before any potential explosive scenario korotayev2020twenty ; korotayev202021st ; modis2013singularity . In this interpretation, the mathematical divergence can be seen as a potential indicator of this deceleration getting closer. In fact, in many systems, it is reasonable to expect a deceleration due to the presence of limiting factors such as the finite availability of resources villani2023super . However, our mathematical analysis reveals that inferring this slowdown from the data may be very tricky. First, the discrete versions of the models, which typically represent the actual dynamics of real-world systems, do not exhibit any singularity. The discrete models can be indeed used to forecast growth beyond this theoretical limit, as detailed in Appendix C. Secondly, our models do not explicitly account for external limiting factors, which are the true drivers of any changes in the system. As a result, these simple models are not able to predict any shift in the system’s evolution. This makes the singularity value practically unrelated to any constraints or deceleration.
@@
Conversely, some authors, particularly in the Big History field, view the singularity as an indicator of upcoming transitions or technological shifts kurzweil2005singularity ; panov2020singularity ; nazaretyan2015megahistory . For instance, the concept of Threshold 9 in the field of Big History nazaretyan2016non suggests that humanity is approaching a profound global paradigm shift (the threshold) that will significantly impact both technology and evolution kurzweil2005singularity . Koppl et al. koppl2023explaining note, however, that the nature of a technological singularity remains uncertain—it ”could be heaven, hell, or something else altogether.” Other authors even speculate about post-singularity civilisations panov2011post . This second interpretation often relies on extrapolations from data without direct evidence that current growth rates will continue indefinitely, positioning such analyses more within the realm of philosophical speculation, though they inspire fascinating and stimulating discussions. From a technical perspective, attempts to extrapolate precise years from the data can lead to peculiar results. For example, in von1960doomsday , the authors calculated the value $N_{t}=1$ from the world population curve and interpreted it as ”Adam’s birth,” resulting in a date nearly 200 billion years in the past. Similarly, some predictions appear strikingly accurate but are essentially coincidental, such as the singularity in GDP per capita growth estimated for 2005 korotayev2009compact , which is closely aligned with the 2008 financial crisis.
@@
Finally, it is important to stress that estimating a reliable singularity value is almost impossible due to the high value of statistical error. As noted in von1960doomsday , the year obtained from fitting data can vary substantially depending on the choice of the last observation. For instance, if Charlemagne, with data until 800 CE, had predicted a singularity, he would have projected it 300 years into the future, while Napoleon, using more recent data, would have predicted it just 30 years ahead. In the very same way, our estimates of singularities, presented in Appendix C, differ from previous analyses in the literature due to the updated data. In certain cases, large time intervals considered in the analysis result in wide confidence bounds for the numerical value of the singularity, making it even more difficult to obtain a reliable estimate.
@@ V Conclusions
This study investigates the dynamics of growth and innovation, offering a unified approach for modelling and understanding it. Innovation dynamics often exhibit hyperbolic trends, which have been studied and interpreted in various ways in the literature. However, a coherent framework that models these processes from a physical and mathematical perspective has been lacking. By leveraging a combinatorial approach based on the Theory of Adjacent Possible (TAP), we provide a comprehensive theoretical foundation to interpret these phenomena. We investigated several distinct scenarios, demonstrating how our framework can successfully replicate the main characteristics of growth trends. We analysed four key cases: biological milestones (Canonical Milestones and Biological Phase Transitions), world population, GDP growth, and technological innovations as seen through U.S. patent data. In each case, the models allow us to estimate the macrodevelopment parameter from the data, enabling us to quantify the system’s growth rate.
This paper demonstrates how a combinatorial framework can offer a coherent interpretation of growth and innovation processes across diverse domains. Future research should explore applying these models to other complex systems, such as biological or evolutionary processes, where similar growth dynamics are observed. The main limitation of our approach is that it does not consider any external factor acting on the system, which is relevant in most cases. Given the simplicity and interpretability of our approach, future work could focus on incorporating external constraints, such as resource limitations, to produce more accurate descriptions and forecasts of system behaviour. Another promising direction is investigating the connection between this framework and the triggering mechanisms described in other discovery processes tria2014dynamics ; di2025dynamics ; bellina2024time . In this view, the emergence of innovations is explained through exploring a network, where each novelty opens up the possibility for further developments. Establishing a connection between these frameworks could further advance our understanding of innovation processes and achieve a comprehensive theory of innovation.